OpenStudy (anonymous):

This question is really throwing me off. Michael throws a rock on Planet X at an angle of 59.4◦ with a speed of 30.2 m/s. It reaches a maximum height of 215 m before falling back down to the ground. (a) What is the acceleration due to gravity on Planet X, gX? Write your answer first in units of m/s2 and then in units of g. (b) How far along the x-direction does the rock travel?

OpenStudy (anonymous):

Please show your work! I'd like to understand how to do it too. Thanks :)

OpenStudy (paxpolaris):

for part (a) do you know a formula including height, initial and final velocities, and acceleration \((-g_X)\)

OpenStudy (paxpolaris):

?

OpenStudy (anonymous):

\[\Delta x = v _{ox} \Delta t\] \[\Delta y=v _{oy} \Delta t + 1/2 a _{y} \Delta t ^{2}\] I have those, but then I'd be missing Delta t, a in y direction and delta x.

OpenStudy (paxpolaris):

something like \[\Large {v}^2={v_0}^2+2a(\Delta x)\]

OpenStudy (anonymous):

This can be used for projectile motion?

OpenStudy (paxpolaris):

sorry it won't be Δx.... why not? \[\Large v^2={v_0}^2+2as \]

OpenStudy (anonymous):

Sorry, what does the s stand for?

OpenStudy (paxpolaris):

s= displacement ... velocites, acceleration, height are all the y components only s=Δy = height

OpenStudy (anonymous):

Alright, thank you! I'll try that out. My prof gave us two equations to use for projectile motion; the ones I stated above. Fingers crossed this works.

OpenStudy (paxpolaris):

\[\Large v=0m/s\] \[\Large v_0=30.2\times\sin\left(59.4\deg\right)m/s\] \[\Large s=215m\] \[\Large a=-g_X\]

OpenStudy (anonymous):

Why is V not 30.2 m/s?

OpenStudy (paxpolaris):



OpenStudy (paxpolaris):

you can switch v and v_0 around, but then \(\Delta y = -215m\)